OpenStudy (dtan5457):

Given the reaction 2C2H6(g)+7O2(g)>>>4CO2(g)+6H2O(g) At STP, what is the total volume of Co2(g) formed when 6.0 liters of C2H6(g) is completely oxidized?

OpenStudy (dtan5457):

@aaronq

OpenStudy (dtan5457):

2 moles of c2h6 is 44.8 liters, so you would divide by 7.46 to get 6 liters... so 4 moles of co2 is 89.6 liters, divided by 7.46=roughly 12 liters? apparently that's the wrong answer, however..

OpenStudy (aaronq):

I was going to say dont worry about scaling down, you can treat the volume just like moles \(\sf \dfrac{V_{ethane}}{2}=\dfrac{V_{CO_2}}{4}\rightarrow \dfrac{6~L}{2}=\dfrac{V_{CO_2}}{4}\) \(\sf V_{CO_2}=12~L\) I got the same answer though

OpenStudy (dtan5457):

the regents says it's 4 liters, but based on the question right before it, which the answer key was wrong. i wouldn't be surprised if this question was wrong too.

OpenStudy (aaronq):

the answer key is probably wrong, often they recycle questions but use different numbers

OpenStudy (dtan5457):

yeah, well thanks for confirming i was right, the book is getting on my nerves lol

OpenStudy (aaronq):

haha that's happened to me too. It's so discouraging when you \(think\) you keep getting the wrong answer